OpenStudy (anonymous):

If the arc on a particular circle has an arc length of 14 inches, and the circumference of the circle is 84 inches, what is the angle measure of the arc

OpenStudy (lieutenantgeneral):

14/84=X(arc length)/360; Multiply both sides by 360=(14x360)/84=X; X=5040/84=60=60 degrees